OpenStudy (anonymous):

Find the area of a parallelogram with sides of 6 and 12 and an angle of 60°. 72√3 sq. units 36√3 sq. units 12√3 sq. units

OpenStudy (anonymous):

The sides of a rhombus with angle of 60° are 6 inches. Find the area of the rhombus. 9√3 in2 18√3 in2 36 in2

OpenStudy (wolf1728):

The area of a parallelogram is equal to the base * height The angle is 60 degrees so the height can be determined by sin(60) = height/6 height = sin(60) * 6 height = 0.86603 * 6 height = 5.19618 So the area = 5.19618 * 6 = 62.35416 (36 * sqrt(3) = 36 * 1.732 = 62.35